# Taming Hyperparameter Sensitivity in Data Attribution: Practical Selection Without Costly Retraining

### v1

@@ B.1 Relationship between TRAK and IFFIM
Because $g_{z^{\prime}}=\frac{1}{n}J^{\top}\alpha_{z^{\prime}}$, for $\mu_{i}$ equal to 00, $\tilde{g}_{z^{\prime},i}$ must also be 00, which can be easily shown through the singular value decomposition of $J$. Hence, B.9 holds when $C_{3}$ is large enough. However, we note here that the sufficient condition for $\dot{c_{p}}(\lambda;z^{\prime})>0$ becomes tighter when $C_{3}$ is larger and $\varepsilon_{g}$ is smaller (see Lemma B.15 for details).
@@
A special case for this assumption is when $J$ has rank $n-1$, where the deducted $1$ rank is because $\sum_{i=1}^{n}\nabla_{\theta}L(z_{i},\theta_{S}^{*})=0$. In this case, since $\sum_{i=1}^{n}\alpha_{z^{\prime},i}=0$ by simple computation, $\alpha_{z^{\prime}}$ is in the column space of $J$.
@@
Assume B.11 holds. For $i\in[n]$, if $\mu_{i}=0$, then the $i$th component of $\alpha_{z^{\prime}}$ under the eigenbasis of $\frac{1}{n}JJ^{\top}$ is also zero.
@@
which is zero for $\mu_{i}=0$ by Lemma B.4. ∎
@@
Under B.7, B.9, and B.11, for all $\lambda>0$,
@@
Denote $w_{z^{\prime},i}$ as the $i$th component of $\alpha_{z^{\prime}}$ under the eigenbasis of $\frac{1}{n}JJ^{\top}$ for $i\in[n]$. We first show that for $i\in[n]$, if $\mu_{i}=0$ then $w_{z^{\prime},i}=0$; if $\mu_{i}\neq 0$, (then $i\leq p$ by definition of $\mu_{i}$) $w_{z^{\prime},i}=\tilde{g}_{z^{\prime},i}\sqrt{n/\mu_{i}}$. For $\mu_{i}=0$, Lemma B.13 guarantees that the component of $\alpha_{z^{\prime}}$ is 00. For $\mu_{i}\neq 0$, with the singular value decomposition of $J$ in Lemma B.4,
@@
Finally, by combining all the results, we state the following formal version of Lemma 4.2.
Lemma B.15 (Formal version of Lemma 4.2).
Assume B.7, B.9, and B.11 hold. With high probability over the sample of $z^{\prime}$ from the test distribution, $r_{z^{\prime},0^{+}}:=\lim_{\lambda\rightarrow 0^{+}}r_{z^{\prime},\lambda}$ exists. Further assume $r_{z^{\prime},0^{+}}>0$. Then if $\mu_{\mathrm{min}}<C_{\mu}$ where $C_{\mu}$ is some positive value depending on $z^{\prime}$, there exists some $C>0$ such that for $0<\lambda<C$,
@@
$C_{\mu}:=(\frac{r_{z^{\prime},0^{+}}(1-p(z^{\prime},\theta_{S}^{*}))\sqrt{C_{1}}}{2C_{2}\sqrt{n\cdot C_{3}}})^{2/\varepsilon_{g}}>0,$
@@
$C:=\min\{C^{*},(\frac{r_{z^{\prime},0^{+}}(1-p(z^{\prime},\theta_{S}^{*}))\sqrt{C_{1}}}{2C_{2}\sqrt{n\cdot C_{3}}})^{2/3}\mu_{\mathrm{min}}^{1-\varepsilon_{g}/3}-\mu_{\mathrm{min}}\}.$
@@
$\mathrm{LHS}\geq\frac{r_{z^{\prime},\lambda}}{\sqrt{n\cdot C_{3}}}\mu_{\mathrm{min}}^{(1-\varepsilon_{g})/2}>\frac{r_{z^{\prime},0^{+}}}{2\sqrt{n\cdot C_{3}}}\mu_{\mathrm{min}}^{(1-\varepsilon_{g})/2}>\frac{C_{2}}{(1-p(z^{\prime},\theta_{S}^{*}))\sqrt{C_{1}}}\frac{(\mu_{\mathrm{min}}+\lambda)^{3/2}}{\mu_{\mathrm{min}}}>\mathrm{RHS},$
@@
Here we show that, in the special case where $z^{\prime}=z_{i}\in S$ for some $i\in[n]$, we have $r_{z^{\prime},0^{+}}>0$. We write the singular value decomposition $J=U\Sigma V^{\top}$. Then by Lemma B.4,
@@
where $e_{i}$ is the $i$th standard basis vector. By Lemma B.13, $\mu_{j}=0$ implies $(U^{\top}\alpha_{z^{\prime}})_{j}=0$. Therefore,
@@
B.4 Effects of Gradient Projection
Our derivation does not rely on specific properties of $J$. When incorporating gradient projection via Eq. 4, as outlined in Remark 4.5, the projection matrix $P$ can be systematically applied to all gradient terms. We now analyze how gradient projection affects $\mu_{\mathrm{min}}$, the smallest non-zero eigenvalue of $F_{S}$ (equivalently, of $\frac{1}{n}JJ^{\top}$. Under projection, this eigenvalue corresponds to the smallest non-zero eigenvalue of $\frac{1}{n}JPP^{\top}J^{\top}$.
In practical applications, gradient projection approximately preserves the dominant eigenvalues of $F_{S}$. Consequently, when $F_{S}$ exhibits vanishingly small (but non-zero) eigenvalues near index $\min\{n,\tilde{p}\}$, the projected matrix $\frac{1}{n}JPP^{\top}J^{\top}$ retains a comparably small non-zero eigenvalue, validating the critical condition in Lemma B.15.
Unlike Arnoldi-based method [32] that aims to preserve top eigenvalues by design, we resort to the standard subspace embedding results [35] for random projection based methods [6, 28]: For common choices of $P$ (e.g. Gaussian random projection) and $10<\varepsilon_{\mathrm{rp}},\delta<$ 1, if $\tilde{p}=\Theta((n+\ln(1/\delta))\varepsilon_{\mathrm{rp}}^{-2})$, then with probability $1-\delta$, $P$ satisfies that for any $v\in\mathbb{R}^{n}$,
@@ B.4 Effects of Gradient Projection
B.5 Proof of Proposition 4.7
Proof of Proposition 4.7.
We utilize Lemma B.3 to obtain
@@
We point out that LHS of Eq. 8 lying in $[0,1]$ is the direct result of applying the generalized Cauchy-Schwarz inequality on $-J\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})$ and $\alpha_{z^{\prime}}$ with inner product defined by positive definite matrix $\frac{1}{n}(\frac{1}{n}JJ^{\top}+\lambda I_{n})^{-1}$, assuming $r_{z^{\prime},\lambda}>0$. Similarly, the RHS of Eq. 8 is shown bounded when the inequality is applied on $-(\frac{1}{n}JJ^{\top}+\lambda I_{n})^{-1}J\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})$ and $-J\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})$ with the same inner product. We use the fact that $t_{2,z^{\prime},\lambda}\geq 0$ as $(\frac{1}{n}JJ^{\top}+\lambda I_{n})^{-2}$ is positive semi-definite. ∎
@@ C.1 Visualization of the Average Surrogate Indicator
To provide better insights about the proposed surrogate indicator, we present the curve of $\bar{\xi}_{T,\lambda}$ as a function of $\lambda$ in Figure 5. In all experiment settings, empirically, $\bar{\xi}_{T,\lambda}$ is a monotonic function of $\lambda$ with a transitional phase near $\bar{\xi}_{T,\lambda}=0.5$, indicating a good sensitivity to $\lambda$ around this value, and supporting our choice of threshold in Algorithm 1.
Figure 5: The curve of $\bar{\xi}_{T,\lambda}$ as a function of $\lambda$. Each subfigure corresponds to an experiment setting outlined in Section 4.3.
@@ C.3 Effects of Different Subset Fractions
We provide an extended analysis of how the subset fraction $a/|S|$ influences the surrogate indicator, considering values in $\{0.25,0.5,0.75\}$, where $0.5$ corresponds to the setting used in earlier experiments. Figure 6 illustrates the surrogate indicator’s behavior for each subset fraction, shown in green ($0.25$), blue ($0.5$), and violet ($0.75$). While the magnitude of the LDS varies with the subset fraction, its overall trend as a function of $\lambda$ remains similar across different values of subset fraction. This consistency suggests that the surrogate indicator is robust to changes in subset size across different experimental settings.

### v2

@@ B.2 LOO Weighted Loss Gradients
Because $g_{z^{\prime}}=\frac{1}{n}J^{\top}\alpha_{z^{\prime}}$, for $\mu_{i}$ equal to 0, $\tilde{g}_{z^{\prime},i}$ must also be 0, which can be easily shown through the singular value decomposition of $J$. Hence, Eq.˜15 holds when $C_{3}$ is large enough. However, we note here that the sufficient condition for $\dot{c_{p}}(\lambda;z^{\prime})>0$ becomes tighter when $C_{3}$ is larger and $\varepsilon_{g}$ is smaller (see Lemma˜B.15 for details).
@@
A special case for this assumption is when $J$ has rank n1n-1, where the deducted $1$ rank is because $\sum_{i=1}^{n}\nabla_{\theta}L(z_{i},\theta_{S}^{*})=0$. In this case, since $\sum_{i=1}^{n}\alpha_{z^{\prime},i}=0$ by simple computation, $\alpha_{z^{\prime}}$ is in the column space of $J$.
@@
Assume Assumption˜B.11 holds. For $i\in[n]$, if $\mu_{i}=0$, then the $i$th component of $\alpha_{z^{\prime}}$ under the eigenbasis of $\frac{1}{n}JJ^{\top}$ is also zero.
@@
which is zero for $\mu_{i}=0$ by Lemma˜B.4. ∎
@@
Under Assumption˜B.7, Eq.˜15, and Assumption˜B.11, for all $\lambda>0$,
@@
Denote $w_{z^{\prime},i}$ as the $i$th component of $\alpha_{z^{\prime}}$ under the eigenbasis of $\frac{1}{n}JJ^{\top}$ for $i\in[n]$. We first show that for $i\in[n]$, if $\mu_{i}=0$ then $w_{z^{\prime},i}=0$; if $\mu_{i}\neq 0$, (then $i\leq p$ by definition of $\mu_{i}$) $w_{z^{\prime},i}=\tilde{g}_{z^{\prime},i}\sqrt{n/\mu_{i}}$. For $\mu_{i}=0$, Lemma˜B.13 guarantees that the component of $\alpha_{z^{\prime}}$ is 0. For $\mu_{i}\neq 0$, with the singular value decomposition of $J$ in Lemma˜B.4,
@@
Finally, by combining all the results, we state the following formal version of Lemma˜4.2.
Lemma B.15 (Formal version of Lemma˜4.2).
Assume Assumption˜B.7, Eq.˜15, and Assumption˜B.11 hold. With high probability over the sample of $z^{\prime}$ from the test distribution, $r_{z^{\prime},0^{+}}\mathrel{\mathop{\ordinarycolon}}=\lim_{\lambda\rightarrow 0^{+}}r_{z^{\prime},\lambda}$ exists. Further assume $r_{z^{\prime},0^{+}}>0$. Then if $\mu_{\mathrm{min}}<C_{\mu}$ where $C_{\mu}$ is some positive value depending on $z^{\prime}$, there exists some $C>0$ such that for $0<\lambda<C$,
@@
$C_{\mu}\mathrel{\mathop{\ordinarycolon}}=(\frac{r_{z^{\prime},0^{+}}(1-p(z^{\prime},\theta_{S}^{*}))\sqrt{C_{1}}}{2nC_{2}\sqrt{n\cdot C_{3}}})^{2/\varepsilon_{g}}>0,$
@@
$C\mathrel{\mathop{\ordinarycolon}}=\min\{C^{*},(\frac{r_{z^{\prime},0^{+}}(1-p(z^{\prime},\theta_{S}^{*}))\sqrt{C_{1}}}{2nC_{2}\sqrt{n\cdot C_{3}}})^{2/3}\mu_{\mathrm{min}}^{1-\varepsilon_{g}/3}-\mu_{\mathrm{min}}\}.$
@@
$\mathrm{LHS}\geq\frac{r_{z^{\prime},\lambda}}{\sqrt{n\cdot C_{3}}}\mu_{\mathrm{min}}^{(1-\varepsilon_{g})/2}>\frac{r_{z^{\prime},0^{+}}}{2\sqrt{n\cdot C_{3}}}\mu_{\mathrm{min}}^{(1-\varepsilon_{g})/2}>\frac{nC_{2}}{(1-p(z^{\prime},\theta_{S}^{*}))\sqrt{C_{1}}}\frac{(\mu_{\mathrm{min}}+\lambda)^{3/2}}{\mu_{\mathrm{min}}}>\mathrm{RHS},$
@@
Here we show that, in the special case where $z^{\prime}=z_{i}\in S$ for some $i\in[n]$, we have $r_{z^{\prime},0^{+}}>0$. We write the singular value decomposition $J=U\Sigma V^{\top}$. Then by Lemma˜B.4,
@@
where $e_{i}$ is the $i$th standard basis vector. By Lemma˜B.13, $\mu_{j}=0$ implies $(U^{\top}\alpha_{z^{\prime}})_{j}=0$. Therefore,
@@
Empirical assessment of constants in assumptions.
The results derived above are dependent on the constants $C_{1}$, $C_{2}$, and $C_{3}$ introduced in Assumptions˜B.7 and 15. Here we provide a brief discussion on the empirical behavior of these constants, showing that they generally stay around the level of $0.1$ to $10$ instead of scaling with the smallest eigenvalue $\mu_{\mathrm{min}}$. Note that, in this empirical study, we compute $C_{1}$ with respect to the smallest eigenvalue rather than all eigenvalues, which is reasonable because the only $C_{1}$ term that occurs in the theoretical derivation is the one related to $\mu_{\mathrm{min}}$. Further, due to computational limitations, we consider projection dimension $4096$. We fix $\varepsilon_{g}=0.5$.
To empirically estimate these constants, we first diagonalize $P^{\top}F_{S}P$ to obtain its eigenvalues and execute a data attribution pass with IFFIM to compute $\nabla_{\theta}L(z^{\prime},\theta^{*}_{S})$ and $g_{z^{\prime}}$ for $z^{\prime}$ in the test dataset. Then, we follow Eq.˜13 and Eq.˜15 to obtain values of $C_{1}$, $C_{2}$, and $C_{3}$ that satisfy these inequalities for all or a given percentage of test examples. For ResNet-9 [14] on CIFAR-2 dataset [24], we find that $C_{3}=0.4082$ is enough for Eq.˜15 to hold on all the test examples, and $C_{1}=0.3121$ and $C_{2}=51.9933$ satisfy Eq.˜13 on $90\%$ of the test examples. In comparison, the smallest eigenvalue is only about $10^{-4}$. For MusicTransformer [2] on MAESTRO dataset [13], we find that $C_{3}=0.0156$ is enough for all test examples, and $90\%$ of test examples yield $C_{1}=0.0518$ and $C_{2}=21.5589$. In comparison, the smallest eigenvalue is only about $7\times 10^{-5}$. These results suggest that, even in highly non-convex settings, our assumptions tend to hold empirically.
B.5 Effects of Gradient Projection
Our derivation does not rely on specific properties of $J$. When incorporating gradient projection via Eq.˜4, as outlined in Remark˜4.5, the projection matrix $P$ can be systematically applied to all gradient terms. We now analyze how gradient projection affects $\mu_{\mathrm{min}}$, the smallest non-zero eigenvalue of $F_{S}$ (equivalently, of $\frac{1}{n}JJ^{\top}$. Under projection, this eigenvalue corresponds to the smallest non-zero eigenvalue of $\frac{1}{n}JPP^{\top}J^{\top}$.
In practical applications, gradient projection approximately preserves the dominant eigenvalues of $F_{S}$. Consequently, when $F_{S}$ exhibits vanishingly small (but non-zero) eigenvalues near index $\min\{n,\tilde{p}\}$, the projected matrix $\frac{1}{n}JPP^{\top}J^{\top}$ retains a comparably small non-zero eigenvalue, validating the critical condition in Lemma˜B.15.
Unlike Arnoldi-based method [33] that aims to preserve top eigenvalues by design, we resort to the standard subspace embedding results [36] for random projection based methods [6, 29]: For common choices of $P$ (e.g. Gaussian random projection) and $0<\varepsilon_{\mathrm{rp}},\delta<1$, if $\tilde{p}=\Theta((n+\ln(1/\delta))\varepsilon_{\mathrm{rp}}^{-2})$, then with probability $1-\delta$, $P$ satisfies that for any $v\in\mathbb{R}^{n}$,
@@ B.5 Effects of Gradient Projection
B.6 Proof of Proposition 4.7
Proof of Proposition˜4.7.
We utilize Lemma˜B.3 to obtain
@@
We point out that LHS of Eq.˜8 lying in $[0,1]$ is the direct result of applying the generalized Cauchy-Schwarz inequality on $J\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})$ and $\alpha_{z^{\prime}}$ with inner product defined by positive definite matrix $\frac{1}{n}(\frac{1}{n}JJ^{\top}+\lambda I_{n})^{-1}$, assuming $r_{z^{\prime},\lambda}>0$. Similarly, the RHS of Eq.˜8 is shown bounded when the inequality is applied on $(\frac{1}{n}JJ^{\top}+\lambda I_{n})^{-1}J\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})$ and $J\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})$ with the same inner product. We use the fact that $t_{2,z^{\prime},\lambda}\geq 0$ as $(\frac{1}{n}JJ^{\top}+\lambda I_{n})^{-2}$ is positive semi-definite. ∎
@@ C.1 Visualization of the Average Surrogate Indicator
To provide better insights about the proposed surrogate indicator, we present the curve of $\bar{\xi}_{T,\lambda}$ as a function of $\lambda$ in Figure˜5. In all experiment settings, empirically, $\bar{\xi}_{T,\lambda}$ is a monotonic function of $\lambda$ with a transitional phase near $\bar{\xi}_{T,\lambda}=0.5$, indicating a good sensitivity to $\lambda$ around this value, and supporting our choice of threshold in Algorithm˜1.
Figure 5: The curve of $\bar{\xi}_{T,\lambda}$ as a function of $\lambda$. Each subfigure corresponds to an experiment setting outlined in Section˜$4.$
@@ C.3 Effects of Different Subset Fractions
We provide an extended analysis of how the subset fraction $a/|S|$ influences the surrogate indicator, considering values in $\{0.25,0.5,0.75\}$, where $0.5$ corresponds to the setting used in earlier experiments. Figure˜6 illustrates the surrogate indicator’s behavior for each subset fraction, shown in green ($0.25$), blue ($0.5$), and violet ($0.75$). While the magnitude of the LDS varies with the subset fraction, its overall trend as a function of $\lambda$ remains similar across different values of subset fraction. This consistency suggests that the surrogate indicator is robust to changes in subset size across different experimental settings.
@@
C.4 Additional Experiments on WikiText2 with GPT2
Figure 7: The plot of LDS versus $\lambda$ on the WikiText2 with GPT2 setting. The left plot shows the results with the IFFIM attributor while the right plot demonstrates the results with the LoGra attributor. The red solid vertical line indicates the $\lambda$ selected by our method.
Here, we demonstrate additional experiments on the large setting WikiText2 [28] with GPT2 [31]. We fix projection dimension $4096$, and apply the surrogate indicator to find proper regularization strengths. We investigate both the IFFIM and the LoGra [6] attributors. The results are illustrated in Figure˜7.
C.5 Experiments on Downstream Settings
We investigate various settings, including Logistic Regression models (LR) on MNIST dataset [25], ResNet-9 [14] on CIFAR-2 [24] dataset, and MusicTransformer (MT) [2] on MAESTRO dataset [13]. We conduct the experiments with training data removal rate $10\%$, $30\%$, and $50\%$. For methods we use for data removal, we have Random (random removal), IFFIM Default and TRAK Default (IFFIM and TRAK attributors with default regularization 0), and IFFIM Selected and TRAK Selected (IFFIM and TRAK attributors with selected regularization by our algorithm). For reference, we also include Full, which stands for test performance without removal. We repeat the experiments with $10$ different seeds and report the mean values as well as their standard errors for better statistical significance. The results are presented in Table˜2, Table˜3, and Table˜4. We notice that when removing $10\%$ of the training data, most methods lead to insignificant model performance decrease, likely due to redundancy in the datasets. However, we observe that our methods (IFFIM Selected and TRAK Selected) generally outperform baselines (Random, IFFIM Default, and TRAK Default) across different settings and the differences are statistically significant.
Table 2: Data selection performance on MNIST+LR settings. Values in the table are means and standard errors of test accuracies. Experiments are repeated with $10$ different seeds. Lower accuracies indicate better attribution performance.
Removal rates 10% 30% 50%
Full $8.63\%\pm 0.03\% 88$$63\%\pm 0.03\% 88.$$8.63\%\pm 0.03\%$
Random $8.52\%\pm 0.03\% 88$$8.06\%\pm 0.05\%$ $87.60\%\pm 0.06\%$
IFFIM Default $88.44\%\pm 0.05\%$ $87.77\%\pm 0.08\%$ $87.58\%\pm 0.10\%$
IFFIM Selected $87.66\%\pm 0.04\%$ $84.50\%\pm 0.04\%$ $81.53\%\pm 0.06\%$
TRAK Default $88.53\%\pm 0.06\%$ $87.92\%\pm 0.11\%$ $86.88\%\pm 0.29\%$
TRAK Selected $\mathbf{87.30\%\pm 0.04\%}$ $\mathbf{83.84\%\pm 0.05\%}$ $\mathbf{80.12\%\pm 0.08\%}$
Table 3: Data selection performance on CIFAR-2+ResNet-9 settings. Values in the table are means and standard errors of test accuracies. Experiments are repeated with $10$ different seeds. Lower accuracies indicate better attribution performance.
Removal rates 10% 30% 50%
Full $75.04\%\pm 0.53\%$ $75.04\%\pm 0.53\%$ $75.04\%\pm 0.53\%$
Random $74.20\%\pm 0.59\%$ $71.86\%\pm 0.33\%$ $68.64\%\pm 0.38\%$
IFFIM Default $74.36\%\pm 0.38\%$ $72.42\%\pm 0.47\%$ $67.84\%\pm 1.16\%$
IFFIM Selected $74.36\%\pm 0.38\%$ $72.42\%\pm 0.47\%$ $67.84\%\pm 1.16\%$
TRAK Default $74.06\%\pm 0.52\%$ $71.68\%\pm 0.40\%$ $69.30\%\pm 0.49\%$
TRAK Selected $\mathbf{71.24\%\pm 0.52\%}$ $\mathbf{64.14\%\pm 0.47\%}$ $\mathbf{56.30\%\pm 0.70\%}$
Table 4: Data selection performance on MAESTRO+MT settings. Values in the table are means and standard errors of test losses. Experiments are repeated with $10$ different seeds. Higher losses indicate better attribution performance. Note that we only include TRAK here, as IF has been shown to exhibit poor attribution performance in this complex setting [8].
Removal rates 10% 30% 50%
Full $4.30\pm 0.01$ $4.30\pm 0.01$ $4.30\pm 0.01$
Random $4.32\pm 0.01$ $4.38\pm 0.01$ $4.67\pm 0.02$
TRAK Default $35\pm 0.01 4.4$$40\pm 0.01$ $4.68\pm 0.01$
TRAK Selected $\mathbf{4.40\pm 0.01}$ $\mathbf{4.52\pm 0.01}$ $\mathbf{4.83\pm 0.02}$
@@ Appendix D Code Availability
Our code is publicly available at https://github.com/TRAIS-Lab/data-attribution-hp.